### v1

# Applications of an identity of Batır

Abstract
Based on an interesting identity of Batır we derive new identities for double sums involving famous number sequences. We also prove some double sum identities for binomial transform pairs.
2010 Mathematics Subject Classification: Primary 05A10, Secondary 11B39, 11B68, 11B83.
Keywords: Double sum, harmonic number, Bernoulli number, Fibonacci number.
1 Introduction
As a reaction to Vălean’s Master Theorem of Series, Batir [4] established a generalization of this theorem. His proof is based on the following identity which holds for any two sequences of complex numbers $(a_{k})_{k\geq 1}$ and $(b_{k})_{k\geq 1}$ [4, Lemma 1]:
$\sum_{k=1}^{n}a_{k}\sum_{j=1}^{k}b_{j}=\sum_{p=0}^{n-1}\sum_{k=1}^{n-p}a_{p+k}b_{k}.$ (1)
Our purpose is to derive various double summation identities based on the following identity which is a variation of (1):
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}a_{n-j}b_{k-j}=\sum_{k=1}^{n}a_{k}\sum_{j=1}^{k}b_{j}.$ (2)
[…]
2 General results
Theorem 1.
If $(a_{k})$, $k=1,2,3,\ldots$, is a sequence of complex numbers, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{a_{n-j}}{n-j}=\sum_{k=1}^{n}a_{k}.$ (3)
Proof.
In (2), set $b_{k}=1$ and replace $a_{k}$ with $a_{k}/k$ for $k=1,2,3,\ldots$
Theorem 2.
If $(a_{k})$, $k=1,2,3,\ldots$, is a sequence of complex numbers, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}(-1)^{k-j+1}a_{n-j}=\sum_{k=0}^{\lfloor(n-1)/2\rfloor}a_{2k+1}.$ (4)
Proof.
In (2), set $b_{k}=(-1)^{k}$ and use
$\sum_{j=1}^{k}(-1)^{j}=\frac{1}{2}\left((-1)^{k}-1\right).$
Proposition 1.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{(n-j)^{2}}=H_{n},$ (5)
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{(n-j)\left(2(n-j)-1\right)}=O_{n}.$ (6)
Proof.
Use $a_{k}=1/k$ and $a_{k}=(2k-1)^{-1}$ in Theorem 1. ∎
Proposition 2.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}(-1)^{k-j+1}(n-j)=\left(\Big\lfloor\frac{n-1}{2}\Big\rfloor+1\right)^{2}.$ (7)
Proof.
Use $a_{k}=k$ in Theorem 2 and
$\sum_{k=0}^{n}(2k+1)=(n+1)^{2}.$
Proposition 3.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}(-1)^{k-j+1}\frac{1}{n-j}=O_{\lfloor(n-1)/2\rfloor}.$ (8)
Proof.
Use $a_{k}=1/k$ in Theorem 2. ∎
Proposition 4.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}(-1)^{k-j+1}F_{n-j}=F_{2(\lfloor(n-1)/2\rfloor+1)}$ (9)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}(-1)^{k-j+1}L_{n-j}=L_{2(\lfloor(n-1)/2\rfloor+1)}-2.$ (10)
Proof.
Use $a_{k}=F_{k}$ and $a_{k}=L_{k}$ in Theorem 2 and combine with
$\sum_{k=0}^{n}F_{2k+1}=F_{2(n+1)}\qquad\text{and}\qquad\sum_{k=0}^{n}L_{2k+1}=L_{2(n+1)}-2.$
Proposition 5.
Let $f(k)$ be any well defined function. If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}(-1)^{k-j}f(n-j)B_{n-j}=\frac{f(1)}{2}.$ (11)
In particular,
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}(-1)^{k-j}B_{n-j}^{2}=-\frac{1}{4}.$ (12)
Proof.
Use $a_{k}=f(k)B_{k}$ in Theorem 2 and the fact that $B_{2k+1}=0$ for $k\geq 1$. ∎
Our next result gives a double sum definition of Bernoulli numbers, expressed in terms of the Stirling numbers of the second kind.
Proposition 6.
If $n$ is a positive integer, then
[…]
Proof.
Use the well-known result (see [8] for annotated references):
[…]
in Theorem 1. ∎
Theorem 3.
If $(a_{k})$, $k=1,2,3,\ldots$, is a sequence of complex numbers, $r$ is a complex number that is not a negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{a_{n-j}}{k-j+r}=\sum_{k=1}^{n}a_{k}H_{k+r}-H_{r}\sum_{k=1}^{n}a_{k}.$ (14)
In particular, we have the following double sum identities
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{a_{n-j}}{k-j}$ $=\sum_{k=1}^{n}{a_{k}H_{k}},$ (15)
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{a_{n-j}}{2k-2j-1}$ $=\sum_{k=1}^{n}{a_{k}O_{k}}.$ (16)
Proof.
Use $b_{j}=1/(j+r)$ in (2). Setting $r=0$ in (14) gives (15) while r=1/2r=-1/2 produces (16). ∎
Proposition 7.
If $n$ is a positive integer, then
[…]
Proof.
Use
[…]
in (15), noting that [10]:
[…]
Proposition 8.
If $n$ is a non-negative integer, $x$ is a complex number with $x\neq 1$, and $r$ is a complex number that is not a negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{x^{n-j}}{k-j+r}=\frac{1}{1-x}\left(\sum_{k=1}^{n}\frac{x^{k}}{k+r}-x^{n+1}(H_{n+r}-H_{r})\right).$ (18)
In particular,
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{x^{n-j}}{k-j}=\frac{1}{1-x}\left(\sum_{k=1}^{n}\frac{x^{k}}{k}-x^{n+1}H_{n}\right)$ (19)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{x^{n-j}}{2k-2j-1}=\frac{1}{1-x}\left(\sum_{k=1}^{n}\frac{x^{k}}{2k-1}-x^{n+1}O_{n}\right).$ (20)
Proof.
Work with $a_{k}=x^{k},x\neq 1,$ in (14). Then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{x^{n-j}}{k-j+r}$ $=\sum_{k=1}^{n}x^{k}H_{k+r}-H_{r}\sum_{k=1}^{n}x^{k}$
$=\sum_{k=1}^{n}x^{k}H_{k+r}-H_{r}\frac{x}{1-x}+H_{r}\frac{x^{n+1}}{1-x}.$
Applying summation by parts we get
$\sum_{k=1}^{n}\frac{x^{k}}{k+r}$ $=\sum_{k=1}^{n}x^{k}(H_{k+r}-H_{k-1+r})$
$=x^{n+1}H_{n+r}-xH_{r}-\sum_{k=1}^{n}H_{k+r}\left(x^{k+1}-x^{k}\right)$
$=x^{n+1}H_{n+r}-xH_{r}+(1-x)\sum_{k=1}^{n}x^{k}H_{k+r}.$
Gathering terms and simplifying proves the main statement. The particular cases are obtained for $r=0$ and r=1/2r=-1/2, respectively. ∎
Remark 1.
The case $x=1$ can be treated with l’Hospital’s rule. However, this case is not interesting in the sense that it is not a double sum as we have for $r\geq 0$ an integer
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{k-j+r}=\sum_{k=1}^{n}\sum_{j=1}^{k}\frac{1}{j+r}=\sum_{k=1}^{n}(H_{k+r}-H_{r}).$
The evaluation of the last sum is elementary.
Corollary 4.
If $n$ is a non-negative integer and $r$ is a complex number that is not a negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{G_{n-j}}{k-j+r}=G_{n+2}(H_{n+r}-H_{r})-\sum_{k=1}^{n}\frac{G_{k+1}}{k+r}.$ (21)
In particular,
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{G_{n-j}}{k-j}=G_{n+2}H_{n}-\sum_{k=1}^{n}\frac{G_{k+1}}{k}$ (22)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{G_{n-j}}{2k-2j-1}=G_{n+2}O_{n}-\sum_{k=1}^{n}\frac{G_{k+1}}{2k-1}.$ (23)
Proof.
Set $x=\alpha$ and $x=\beta$ with $\alpha=(1+\sqrt{5})/2$ and $\beta=-1/\alpha$ in (18), simplify, and combine according to the Binet forms for Fibonacci (Lucas) numbers. ∎
Corollary 5.
If $n$ is a non-negative integer, then
[…]
Proof.
When x=1x=-1 in (19) , then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{(-1)^{n-j}}{k-j}$ $=\frac{1}{2}\left(\sum_{k=1}^{n}\frac{(-1)^{k}}{k}+(-1)^{n}H_{n}\right)$
$=\frac{1}{2}\left(-H_{n}+H_{\lfloor n/2\rfloor}+(-1)^{n}H_{n}\right)$
[…]
Proposition 9.
If $n$ is a non-negative integer and $p$ and $x$ are complex numbers, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{x^{n-j}}{k-j}\left(\frac{1}{n+p-j}-\frac{x}{n+1+p-j}\right)=\sum_{k=1}^{n}\frac{x^{k}}{k(k+p)}-H_{n}\frac{x^{n+1}}{n+1+p}.$ (25)
Proof.
Multiply (19) by $(1-x)x^{p-1},p\in\mathbb{C},$ and integrate. ∎
Corollary 6.
If $n$ is a non-negative integer and $p$ is a complex number that is not a non-positive integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{(k-j)(n+p-j)(n+1+p-j)}=\frac{1}{p}\left(H_{n}+H_{p}-H_{n+p}\right)-\frac{H_{n}}{n+1+p}.$ (26)
In particular,
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{(k-j)(n+1-j)(n+2-j)}=\frac{n}{n+1}-\frac{H_{n}}{n+2}$ (27)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{(k-j)(2(n-j)-1)(2(n-j)+1)}=O_{n}-\frac{n+1}{2n+1}H_{n}.$ (28)
Proof.
Set $x=1$ in (25) and calculate
$\sum_{k=1}^{n}\frac{1}{k(k+p)}=\frac{1}{p}\sum_{k=1}^{n}\left(\frac{1}{k}-\frac{1}{k+p}\right)=\frac{1}{p}\left(H_{n}+H_{p}-H_{n+p}\right).$
Evaluate (26) at $p=1$ and p=1/2p=-1/2, respectively. ∎
Proposition 10.
If $r$ is a non-negative integer, then
$\sum_{m=1}^{r}{\sum_{n=1}^{m-1}{\sum_{k=1}^{n-1}{\sum_{j=0}^{k-1}{\frac{1}{{\left({k-j}\right)\left({n-j}\right)\left({n-j+1}\right)}}}}}}=r-H_{r}-\frac{{r+1}}{2}\left({H_{r}^{2}-H_{r}^{(2)}}\right).$ (29)
Proof.
The sums
$\sum_{m=1}^{r}{H_{m}}$ $=\left({r+1}\right)H_{r}-r,$ (30)
$\sum_{m=1}^{r}H_{m}^{2}$ $=(r+1)H_{r}^{2}-(2r+1)H_{r}+2r,$ (31)
and
$\sum_{m=1}^{r}{H_{m}^{(2)}}=\left({r+1}\right)H_{r}^{(2)}-H_{r}$ (32)
can be derived easily. Write n1n-1 for $n$ in (27) and sum to obtain
$\sum_{n=1}^{m-1}{\sum_{k=1}^{n-1}{\sum_{j=0}^{k-1}{\frac{1}{{\left({k-j}\right)\left({n-j}\right)\left({n-j+1}\right)}}}}}=m-H_{m}-\frac{1}{2}\left({H_{m}^{2}-H_{m}^{(2)}}\right).$ (33)
Summing (33) over $m$ using (30), (31) and (32) produces (29).
Proposition 11.
If $n$ is a non-negative integer and $p$ is a positive integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{(k-j)(n-j+p)}$ $=\frac{1}{2}\left(H_{n}^{2}-H_{n}^{(2)}\right)-\frac{1}{2}\left(H_{p-1}^{2}+H_{p-1}^{(2)}\right)$
$\qquad+H_{p-1}\left(H_{p+n-1}-H_{n}\right)+H_{n}\left(H_{p+n}-H_{n}\right)$
$\qquad\qquad-\sum_{k=1}^{p-1}\frac{H_{k-1}}{n+k}.$ (34)
In particular,
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{(k-j)(n+1-j)}=\frac{1}{2}\left(H_{n}^{2}-H_{n}^{(2)}\right)+\frac{H_{n}}{n+1}.$ (35)
Proof.
Work with $a_{k}=1/(k+p),p\geq 1,$ in (15). The particular case follows by setting $p=1$. ∎
Proposition 12.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{H_{n-j}}{k-j}=(n+1)H_{n}^{2}-(2n+1)H_{n}+2n,$ (36)
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{H_{n-j}^{2}}{k-j}=(n+1)H_{n}^{3}-\frac{3}{2}(2n+1)H_{n}^{2}+3(2n+1)H_{n}+\frac{1}{2}H_{n}^{(2)}-6n,$ (37)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{H_{n-j}^{(2)}}{k-j}=(n+1)H_{n}H_{n}^{(2)}-\frac{1}{2}(2n+1)H_{n}^{(2)}+H_{n}-\frac{1}{2}H_{n}^{2}.$ (38)
Proof.
Work with $a_{k}=H_{k}$, $a_{k}=H_{k}^{2}$, and $a_{k}=H_{k}^{(2)}$, in turn, in (15). The sums
$\sum_{k=1}^{n}H_{k}^{3}=(n+1)H_{n}^{3}-\frac{3}{2}(2n+1)H_{n}^{2}+3(2n+1)H_{n}+\frac{1}{2}H_{n}^{(2)}-6n$
and
$\sum_{k=1}^{n}H_{k}H_{k}^{(2)}=(n+1)H_{n}H_{n}^{(2)}-\frac{1}{2}(2n+1)H_{n}^{(2)}+H_{n}-\frac{1}{2}H_{n}^{2}$
come from the paper [9]. ∎
3 Miscellaneous results involving binomial coefficients
Proposition 13.
If $m$ and $n$ are non-negative integers and $p$ is a complex number, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{(-1)^{j}}{k-j}\binom{p}{j}\binom{n-j}{m}=\binom{n-p}{m-p}\left(H_{n-p}-H_{m-p}+H_{m}\right).$ (39)
Proof.
Use
$a_{k}=(-1)^{k}\binom{p}{n-k}\binom{k}{m}$
in (15), since ([12, Corollary 5])
$\sum_{k=1}^{n}(-1)^{k}\binom{p}{n-k}\binom{k}{m}H_{k}=(-1)^{n}\binom{{n-p}}{{m-p}}\left(H_{n-p}-H_{m-p}+H_{m}\right).$
Proposition 14.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\binom{n}{j}\frac{x^{n-j}}{k-j}=(1+x)^{n}H_{n}-\sum_{k=1}^{n}\frac{(1+x)^{n-k}}{k}.$ (40)
In particular,
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{\binom{n}{j}}{k-j}=2^{n}\left(H_{n}-\sum_{k=1}^{n}\frac{1}{2^{k}k}\right).$ (41)
Proof.
Work with $a_{k}=\binom{n}{k}x^{k}$ in (15) while making use of the Knuth-Boyadzhiev identity
$\sum_{k=1}^{n}\binom{n}{k}H_{k}x^{k}=(1+x)^{n}H_{n}-\sum_{k=1}^{n}\frac{(1+x)^{n-k}}{k}.$ (42)
Proposition 15.
If $m$ and $n$ are non-negative integers, then
[…]
Proof.
Write $\exp(x)$ for $x$ in (40) and differentiate $m$ times through the resulting equation and evaluate at $x=0$, using
[…]
Proposition 16.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{\binom{n}{j}^{2}}{k-j}=\binom{{2n}}{n}\left(2H_{n}-H_{2n}\right)$ (44)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{k-j}\binom{n}{j}\binom{2n}{n+j}=\binom{3n}{n}H_{n}-\sum_{k=1}^{n}\frac{1}{k}\binom{3n-k}{n-k}$. (45)
Proof.
To derive Identity (44), plug $a_{k}=\binom{n}{k}^{2}$ in (15) and use the fact that [6]:
$\sum_{k=0}^{n}{\binom{n}{k}^{2}H_{k}}=\binom{{2n}}{n}\left(2H_{n}-H_{2n}\right).$
Identity (45) is obtained by setting $a_{k}=\binom{n}{k}\binom{2n}{k}$ in (15) and making use of [11]:
$\sum_{k=0}^{n}{\binom{n}{k}\binom{2n}{k}H_{k}}=\binom{3n}{n}H_{n}-\sum_{k=1}^{n}\frac{1}{k}\binom{3n-k}{n-k}$.
Proposition 17.
If $n$ is a non-negative integer, then
[…]
Proof.
Set $a_{k}=\binom{n}{k}^{3}$ in (3) and use Dixon’s identity [7]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}^{3}}=\begin{cases}(-1)^{n/2}\binom{{n}}{{n/2}}\binom{{3n/2}}{n},&\text{if $n$ is even;}\\ 0,&\text{if $n$ is odd.}\end{cases}$
Proposition 18.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{2^{2j}}{k-j}\binom{2(n-j)}{n-j}=2^{2n+1}+\left(H_{n}-2\right)(2n+1)\binom{2n}{n}$ (47)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{2^{2j}}{2k-2j-1}\binom{2(n-j)}{n-j}=\left(O_{n+1}-1\right)(2n+1)\binom{2n}{n}$. (48)
Proof.
Use $a_{k}=2^{-2k}\binom{2k}{k}$ in (15) and also in (16) and employ the following identities from [3]:
$\sum_{k=1}^{n}2^{-2k}\binom{2k}{k}H_{k}=2+\left(H_{n}-2\right)(2n+1)2^{-2n}\binom{{2n}}{n}$,
$\sum_{k=1}^{n}2^{-2k}\binom{2k}{k}O_{k}=\left(O_{n+1}-1\right)(2n+1)2^{-2n}\binom{{2n}}{n}$.
Proposition 19.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{2^{2j}}{k-j}\binom{2(n-j)}{n-j}O_{n-j}=\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{2^{2j}}{2k-2j-1}\binom{2(n-j)}{n-j}H_{n-j}.$ (49)
Proof.
This result follows from (15) and (16) and the following identity [5]:
$\sum_{k=1}^{n}2^{-2k}\binom{2k}{k}O_{k}H_{k}$ $=(2n+1)2^{-2n}\binom{2n}{n}\left(H_{n}-2\right)\left(O_{n}-\frac{2n}{2n+1}\right)$
$\qquad+\frac{2n+1}{2(4n-1)}\binom{2n}{n}-2.$
Proposition 20.
If $n$ is a non-negative integer and $s$ is a complex number that is not a negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{(-1)^{j}s}{k-j}\binom{s}{n-j}=s\binom{s-1}{n}H_{n}+\binom{s-1}{n}-(-1)^{n}.$ (50)
Proof.
Batir derived the identity [5, Corollary 3]
$\sum_{k=1}^{n}\binom{n+s}{k}(-1)^{k}H_{k}=(-1)^{n}\binom{n+s-1}{n}H_{n}+\frac{(-1)^{n}}{n+s}\binom{n+s-1}{n}-\frac{1}{n+s}.$
Using the transformation $s\mapsto s-n$ this identity can be stated equivalently as
$\sum_{k=1}^{n}\binom{s}{k}(-1)^{k}sH_{k}=s(-1)^{n}\binom{s-1}{n}H_{n}+(-1)^{n}\binom{s-1}{n}-\frac{1}{s}.$
Now, set $a_{k}=(-1)^{k}\binom{s}{k}$ in (15). ∎
Corollary 7.
If $n$ is a non-negative integer then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{1}{k-j}\frac{2^{2j}}{2(n-j)-1}\binom{2(n-j)}{n-j}=2^{2n+1}-\binom{2n}{n}(H_{n}+2)$ (51)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{2^{2j}}{k-j}\binom{2(n-j)}{n-j}=2^{2n+1}+(2n+1)\binom{2n}{n}(H_{n}-2).$ (52)
Proof.
Evaluate (50) at s=1/2s=-1/2 and $s=1/2$, and keep in mind that
$\binom{1/2}{r}=(-1)^{r+1}\frac{2^{-2r}}{2r-1}\binom{2r}{r},\quad\binom{-1/2}{r}=(-1)^{r}2^{-2r}\binom{2r}{r}$,
and
$\binom{-3/2}{r}=(-1)^{r}(2r+1)2^{-2r}\binom{2r}{r}$.
Proposition 21.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{\frac{{(-1)^{j}}}{{k-j}}\binom{{n}}{j}^{-1}}}=\left({n+1}\right)\left({\frac{{H_{n+1}}}{{n+2}}-\frac{{(-1)^{n}+1}}{{\left({n+2}\right)^{2}}}}\right).$ (53)
Proof.
Use $a_{k}=(-1)^{k}\binom{n}{k}^{-1}$ in (15) together with the following identity [5]:
$\sum_{k=1}^{n}{(-1)^{k}\binom{{n}}{k}^{-1}H_{k}}=\left({n+1}\right)\left({\frac{{H_{n+1}}}{{n+2}}-\frac{{(-1)^{n}+1}}{{\left({n+2}\right)^{2}}}}\right).$
Lemma 1.
If $n$ is a non-negative integer and $p$ and $q$ are complex numbers that are not negative integers, then
$\sum_{k=1}^{n}{\binom{{k+q}}{p}}=\frac{{n+q+1}}{{p+1}}\binom{{n+q}}{p}-\binom{{q}}{{p+1}}-\binom{q}{p}$ (54)
and
$\sum_{k=1}^{n}{\binom{{k+q}}{p}H_{k+q}}$ $=\binom{{n+q}}{p}\left({\frac{1}{{p+1}}+\frac{{n+q+1}}{{p+1}}\left({H_{n+q}-\frac{1}{{p+1}}}\right)}\right)$
$\qquad-\binom{{q}}{{p+1}}\left({H_{q}-\frac{1}{{p+1}}}\right)-\binom{{q}}{p}H_{q}.$ (55)
Proof.
Identity (54) is a consequence of Pascal’s formula:
$\binom{{k+q}}{p}=\binom{{k+q+1}}{{p+1}}-\binom{{k+q}}{{p+1}}$.
Identity (1) is obtained by differentiating (54) first with respect to $p$, and then with respect to $q$ and eliminating $H_{q-p-1}$ between the two resulting equations. ∎
Lemma 2.
If $k$ is a non-negative integer and $r$ is a complex number that is not a negative integer, then
$\binom{{k-1/2}}{k+r}=(-1)^{r}r\binom{{2r}}{r}\frac{{C_{k}}}{{2^{2k+2r}}}\binom{{k+r}}{{k+1}}^{-1},$ (56)
where $C_{j}$ is the $j^{th}$ Catalan number, defined for every non-negative integer $j$ by
$C_{j}=\frac{1}{j+1}\binom{2j}{j}$.
Proof.
A consequence of the generalized binomial coefficient, expressed in terms of the Gamma function, $\Gamma(z)$, namely,
$\binom{r}{s}=\frac{{\Gamma(r+1)}}{{\Gamma(s+1)\Gamma(r-s+1)}}.$
Proposition 22.
If $n$ is a non-negative integer and $p$ and $q$ are complex numbers that are not negative integers, then
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{\frac{1}{{n-j}}\binom{{n-j+q}}{p}}}=\frac{{n+q+1}}{{p+1}}\binom{{n+q}}{p}-\binom{{q}}{{p+1}}-\binom{{q}}{p}$ (57)
and
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{\frac{{\binom{{n-j+q}}{p}}}{{k-j+q}}}}$ $=\binom{{n+q}}{p}\left({\frac{1}{{p+1}}+\frac{{n+q+1}}{{p+1}}\left({H_{n+q}-H_{q}-\frac{1}{{p+1}}}\right)}\right)$
$\qquad+\binom{{q}}{{p+1}}\frac{1}{{p+1}}.$ (58)
Proof.
Identity (57) is obtained by putting $a_{k}=\binom{k+q}{p}$ in Theorem 1 and using (54). Set $a_{k}=\binom{k+q}{p}$ in (14) with $r=q$ and make use of (54) and (1) to get (22). ∎
Our next result is a double sum identity that involves binomial coefficients, Catalan numbers, and odd harmonic numbers.
Proposition 23.
If $n$ is a non-negative integer and $r\neq 1/2$ is a complex number that is not a non-positive integer, then
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{2^{2j}\binom{{n-j+r}}{{r-1}}^{-1}\frac{{C_{n-j}}}{{2k-2j-1}}}}$
$\qquad=-\binom{{n+r}}{{r-1}}^{-1}\frac{{C_{n}}}{{2r-1}}\left({1+\left({2n+1}\right)\left({O_{n}+\frac{1}{{2r-1}}}\right)}\right)+\frac{{2^{2n+1}}}{{\left({2r-1}\right)^{2}}}.$ (59)
In particular,
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{\frac{{2^{2j}C_{n-j}}}{{2k-2j-1}}}}=-C_{n}\left({1+\left({2n+1}\right)\left({O_{n}+1}\right)}\right)+2^{2n+1}.$ (60)
Proof.
With Lemma 2 in mind, set q=1/2q=-1/2 and p=r1/2p=-r-1/2 in (22). ∎
Lemma 3 ([2, Section 13]).
If $n$ is a non-negative integer and $r$ is a complex number that is not a negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{{k+r}}}=\begin{cases}\dfrac{{\sin(\pi r)}}{{\pi\left({n+r}\right)}},&\text{if $r\neq 0$;}\\ 0,&\text{if $r=0$, $n\neq 0$;}\end{cases}$ (61)
and
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{k}}{{k+r}}H_{k+r}}=\begin{cases}\dfrac{{\sin(\pi r)}}{{\pi\left({n+r}\right)}}\left({H_{r-1}+\dfrac{1}{{n+r}}}\right),&\text{if $r\neq 0$;}\\ -\dfrac{1}{n},&\text{if $r=0$, $n\neq 0$.}\end{cases}$ (62)
Proposition 24.
If $n$ is a non-negative integer and $r$ is a complex number that is not a negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{(-1)^{j}}{k-j+r}\binom{n+r}{j}$
$\qquad=(-1)^{n}\binom{n}{n+r}^{-1}\begin{cases}\dfrac{{\sin(\pi r)}}{{\pi\left({n+r}\right)}}\left({\dfrac{1}{n+r}-\dfrac{1}{r}}\right),&\text{if $r\neq 0$;}\\ -\dfrac{1}{n},&\text{if $r=0$, $n\neq 0$.}\end{cases}$ (63)
Proof.
Set $a_{k}=(-1)^{k}\binom{{n}}{k}\binom{{k}}{{k+r}}$ in (14) and use (61) and (62). ∎
Corollary 8.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{(-1)^{j}}{2(k-j)+1}\frac{\binom{2n+1}{2j}\binom{2j}{j}}{\binom{n}{j}}2^{-2j}=(-1)^{n+1}\frac{n(n+1)}{(2n+1)^{2}}2^{-2n}\binom{2(n+1)}{n+1}$. (64)
Proof.
Set $r=1/2$ in (24) and simplify using
$\binom{n+1/2}{j}=\frac{\binom{2n+1}{2j}\binom{2j}{j}}{\binom{n}{j}}2^{-2j}\quad\text{and}\quad\binom{n}{n+1/2}=\frac{1}{\pi}\frac{2^{2n+2}}{n+1}\binom{2(n+1)}{n+1}^{-1}.$
4 Applications to binomial transforms
Two sequences of complex numbers $(s_{n})$ and $(\sigma_{n})$ are called a binomial transform pair if these sequences are connected by the relations
$\sigma_{n}=\sum_{k=0}^{n}\binom{n}{k}(-1)^{k}s_{k}\qquad\iff\qquad s_{n}=\sum_{k=0}^{n}\binom{n}{k}(-1)^{k}\sigma_{k}.$
Theorem 9.
Let $(a_{k})$ be a sequence of complex numbers. Let $(s_{k})$ and $(\sigma_{k})$ be a binomial transform pair. Then
$\sum_{k=1}^{n}(-1)^{k}\sum_{j=0}^{k-1}(-1)^{j}\binom{k}{j}a_{n-j}s_{k-j}=\sum_{k=1}^{n}a_{k}\sigma_{k}-s_{0}\sum_{k=1}^{n}a_{k}.$ (65)
Proof.
Work with (2) and apply it to $b_{j}=\binom{k}{j}(-1)^{j}s_{j}$. Note that
$\sum_{j=1}^{k}b_{j}=\sum_{j=1}^{k}\binom{k}{j}(-1)^{j}s_{j}=\sum_{j=0}^{k}\binom{k}{j}(-1)^{j}s_{j}-s_{0}=\sigma_{k}-s_{0}.$
Proposition 25.
Let $(a_{k})$ be a sequence of complex numbers. Let $F_{n}$ be the Fibonacci sequence and $L_{n}$ be the Lucas sequence, respectively. Then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\binom{k}{j}a_{n-j}F_{k-j}=\sum_{k=1}^{n}a_{k}F_{2k}$ (66)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\binom{k}{j}a_{n-j}L_{k-j}=\sum_{k=1}^{n}a_{k}(L_{2k}-2).$ (67)
Proof.
Apply Theorem 9 in turn to the binomial transform pairs $\{s_{n},\sigma_{n}\}=\{(-1)^{n}F_{n},F_{2n}\}$ and $\{s_{n},\sigma_{n}\}=\{(-1)^{n}L_{n},L_{2n}\}$. ∎
Proposition 26.
Let $(a_{k})$ be a sequence of complex numbers. Let $B_{n}$ be the sequence Bernoulli numbers. Then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\binom{k}{j}a_{n-j}B_{k-j}=\sum_{k=1}^{n}a_{k}\left((-1)^{k}B_{k}-1\right).$ (68)
Proof.
Apply Theorem 9 to the binomial transform pair $\{s_{n},\sigma_{n}\}=\{(-1)^{n}B_{n},(-1)^{n}B_{n}\}$. ∎
Theorem 10.
Let $(s_{k})$ and $(\sigma_{k})$ be a binomial transform pair. Let $m$, $n$, and $r$ be non-negative integers. Then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{(-1)^{j}}{n-j+1}\binom{n}{j}\sum_{p=0}^{r}(-1)^{p}\binom{r}{p}s_{k+p+m-j}$
$\qquad=\frac{(-1)^{n}}{n+1}\sum_{p=0}^{m}(-1)^{p}\binom{m}{p}\left(\sigma_{n+p+r}-\sigma_{p+r}\right).$ (69)
In particular,
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{\frac{{(-1)^{j}}}{{n-j+1}}\binom{{n}}{j}\sum_{p=0}^{r}{(-1)^{p}\binom{{r}}{p}s_{k+p-j}}}}$ $=\frac{{(-1)^{n}}}{{n+1}}\left({\sigma_{n+r}-\sigma_{r}}\right),$ (70)
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{\frac{{(-1)^{j}}}{{n-j+1}}\binom{{n}}{j}s_{k+m-j}}}$ $=\frac{{(-1)^{n}}}{{n+1}}\sum_{p=0}^{m}{(-1)^{p}\binom{{m}}{p}\left({\sigma_{n+p}-\sigma_{p}}\right)},$ (71)
and
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{(-1)^{j}}{n-j+1}\binom{n}{j}s_{k-j}=\frac{(-1)^{n}}{n+1}(\sigma_{n}-\sigma_{0}).$ (72)
Proof.
Set $a_{k}=(-1)^{k}\binom{n}{k}(k+1)^{-1}$ in (2) and apply [2, Theorems 6.3 and 7.9]. ∎
Proposition 27.
If $n$ is a non-negative integer and $x$ is a complex number, then
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{(-1)^{j}}{n-j+1}\binom{n}{j}x^{k-j}H_{k-j}=\frac{(-1)^{n}}{n+1}\left((1-x)^{n}H_{n}-\sum_{k=1}^{n}\frac{(1-x)^{n-k}}{k}\right).$ (73)
In particular,
$\sum_{k=1}^{n}\sum_{j=0}^{k-1}\frac{(-1)^{j}}{n-j+1}\binom{n}{j}H_{k-j}=\frac{(-1)^{n+1}}{n(n+1)}.$ (74)
Proof.
Work in (72) with the binomial transform pair
$s_{n}=x^{n}H_{n}\quad\text{and}\quad\sigma_{n}=(1-x)^{n}H_{n}-\sum_{k=1}^{n}\frac{(1-x)^{n-k}}{k},$
which follows from the Knuth-Boyadzhiev identity (42). ∎
Proposition 28.
If $n$ is a positive integer and $r$ is a non-negative integer such that $r<n$, then
[…]
Proof.
Write $\exp x$ for $x$ in (73) and proceed as in the proof of Proposition 15. ∎
Proposition 29.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{(-1)^{j}2^{2j-2k}\frac{{2(k-j)+1}}{{(n-j+1)(k-j+1)}}\binom{{n}}{j}C_{k-j}\left({H_{k-j+1}-O_{k-j+1}}\right)}}$
$\qquad=(-1)^{n+1}2^{-2n}\frac{{2n+1}}{{\left({n+1}\right)^{2}}}C_{n}\left({H_{n+1}-O_{n+1}}\right).$ (76)
Proof.
Use the following known result ([2, Prop. 13.22]):
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}\frac{{2k+1}}{{k+1}}C_{k}\left({H_{k+1}-O_{k+1}}\right)}=-2^{-2n}\frac{{2n+1}}{{n+1}}C_{n}\left({H_{n+1}-O_{n+1}}\right)$
in (72). ∎
Proposition 30.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}{(-1)^{k}\sum_{j=0}^{k-1}{\frac{{B_{k-j+1}}}{{\left({k-j+1}\right)\left({n-j+1}\right)}}\binom{{n}}{j}}}=\frac{{(-1)^{n}}}{{\left({n+1}\right)^{2}}}\left({(-1)^{n+1}B_{n+1}+\frac{{n-1}}{2}}\right).$ (77)
Proof.
Use the identity ([2, Prop. 13.33]):
$\sum_{k=0}^{n}{\binom{{n}}{k}\frac{{B_{k+1}}}{{k+1}}}=(-1)^{n+1}\frac{{B_{n+1}}}{{n+1}}-\frac{1}{{n+1}}$
in (72). ∎
Proposition 31.
If $n$ is a non-negative integer, then
$\sum_{k=1}^{n}{\sum_{j=0}^{k-1}{\frac{{(-1)^{j}}}{{n-j+1}}\binom{{n}}{j}\binom{{2\left({k-j}\right)}}{{k-j}}}}=\frac{1}{{n+1}}\left({\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{k}\binom{{n-k}}{k}}-(-1)^{n}}\right).$ (78)
Proof.
It is known that [1, Prop. 6]:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\binom{{2k}}{k}}=(-1)^{n}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{k}\binom{{n-k}}{k}}$.
Use this result in (72). ∎